OpenStudy (anonymous):

PLEASE HELP solve the equation and write the solution set: |4x-8|-3x=6x+2

OpenStudy (jim766):

start by getting absolute value signs by themself, how you going to get rid of the -3x?

OpenStudy (anonymous):

im noy sure please explain to me how to do this

OpenStudy (jim766):

To get rid of the -3x, add 3x to both sides..... so you end up with \[\left| 4x-8 \right|= 9x +2\] do you see where that came from?

OpenStudy (jim766):

now solve 2 seperate equations 4x - 8 = 9x +2 and 4x -8 = -(9x + 2)